# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ Abstract
We show that Hall conductance and its non-abelian and higher-dimensional analogs are obstructions to promoting a symmetry of a state to a gauge symmetry. To do this, we define a local Lie algebra over a Grothendieck site as a pre-cosheaf of Lie algebras with additional properties and propose that a gauge symmetry should be described by such an object. We show that infinitesimal symmetries of a gapped state of a quantum spin system form a local Lie algebra over a site of semilinear sets and use it to construct topological invariants of the state. Our construction applies to lattice systems on arbitrary asymptotically conical subsets of a Euclidean space including those which cannot be studied using field theory.
@@ 1 Introduction
The study of gapped phases of quantum matter at zero temperature is an important area of theoretical physics. Much conceptual progress has been made by assuming that gapped phases can be described by topological quantum field theory (TQFT). For example, the celebrated Quantum Hall Effect is captured by Chern-Simons field theory. However, the precise relation between gapped phases of matter and TQFTs is not understood. Recently, new mathematically rigorous approaches to classifying gapped phases of matter have been developed (see [1, 2, 3, 4] for the case of one-dimensional systems, [5, 6, 7, 8, 9, 10] for the case of two-dimensional systems, and [11] for systems in an arbitrary number of dimensions). They enable one to assign indices to gapped states of infinite-volume quantum systems invariant under symmetries. The main property of these indices, also referred to as topological invariants, is that they do not vary along suitably-defined continuous paths in the space of states. In some cases, the indices can be related to physical quantities, such as the zero-temperature Hall conductance, thereby explaining the robustness of the latter.
The methods of [7, 11, 8, 9] apply to arbitrary gapped states of infinite-volume quantum spin systems with rapidly decaying interactions and employ $C^{*}$-algebraic techniques, some well-established and some relatively new. The construction of topological invariants in [11, 8] also uses some algebraic and geometric ingredients . The algebraic ingredient is a pointed (or curved) Differential Graded Lie Algebra (DGLA) and an associated Maurer-Cartan equation. The geometric ingredient is a collection of conical subsets of the Euclidean space triangulating the sphere at infinity. The appearance of these ingredients in the context of quantum statistical mechanics has not been motivated, and consequently the mathematical meaning of the invariants remains obscure.
@@
The primary goal of this paper is provide a proper mathematical framework for the constructions of [11, 8] and to interpret topological invariants of gapped states as lattice analogs of ’t Hooft anomalies in Quantum Field Theory. The secondary goal is to generalize the construction in various directions. In particular, we show how to define topological invariants of lattice spin systems confined to well-behaved subsets of the lattice. This generalization makes explicit that the invariants take values in a vector space which is determined by the asymptotic geometry of the subset.
While our work concerns quantum lattice systems, we take inspiration from Quantum Field Theory (QFT). These two subjects are connected via the bulk-boundary correspondence. One aspect of this conjectural correspondence is that topological invariants of gapped states with symmetries are related to ’t Hooft anomalies of symmetries of the boundary field theory. [^1] This assumes, of course, that a field-theoretic description of boundary degrees of freedom exists, which is far from obvious. It is usually said that ’t Hooft anomalies are obstructions to gauging a global symmetry of a QFT [12]. A possible mathematical interpretation of this statement is that an ’t Hooft anomaly is an obstruction to defining a local action of the group of gauge transformations on the algebra of local observables of a QFT. Assuming this interpretation, the presence of an ’t Hooft anomaly is a purely kinematic statement which involves neither the Hamiltonian nor the vacuum state of the field theory. It is not clear if conventional markers of ’t Hooft anomalies, such as anomalous Ward identities for vacuum correlators of currents, are implied by a kinematic statement. Proving or disproving this is currently out of reach because of gaps in the mathematical foundations of QFT. The mathematical theory of quantum lattice systems, on the other hand, is sufficiently mature and enables us to address the problem of ’t Hooft anomalies from the bulk side of the bulk-boundary correspondence. In this paper we show that topological invariants of gapped states of lattice systems, such as the zero-temperature Hall conductance, can be interpreted as obstructions to promoting a symmetry of a gapped state to a gauge symmetry. Dynamics enters this statement only though the state.
The main novelty of the paper is a new formulation of locality on a lattice. Building on ideas introduced in [7, 11, 8], we define for any (possibly unbounded) region of the lattice a space of derivations that are approximately localized on that region. For sufficiently regular regions, we show that these spaces behave as expected under natural operations like the commutator. To combine them into a single geometric object we show that they form a cosheaf on a certain site, i. e. a category with a Grothendieck topology. The utility of Grothendieck topologies in describing spaces of functions with a prescribed asymptotic behavior is well-known to analysts, see e.g. [13, 14], and here we apply the same idea in a non-commutative context. We call the resulting global geometric object a local Lie algebra. An on-site action of a Lie group on a lattice system yields a representation of a certain local Lie algebra by derivations, and we show that the invariants defined in [11] can be phrased as purely algebraic invariants of this representation. Namely, an obstruction exists to defining a representation which acts by state-preserving derivations, and this obstruction takes value in the homology of a certain DGLA associated naturally to the local Lie algebra of state-preserving derivations. The invariants defined in [11] then arise from a natural pairing of this homology with the Čech cohomology of the sphere at infinity. Aside from clarifying their mathematical meaning, this also shows that the invariants defined in [11] do not depend on certain choices present in their construction.
@@
The content of the paper is as follows. In Section 2 we axiomatize the notion of an infinitesimal local symmetry by defining local Lie algebras abstractly in terms of cosheaves on a site. We also introduce the Čech functor which assigns a DGLAs to a local Lie algebra, and will serve as the main link between lattice geometry and algebraic topology. In Section 3 we turn to infinitesimal symmetries on the lattice, introducing the space of derivations approximately localized on a region, and prove the main properties of these spaces, some of which hold only for sufficiently regular regions. In section 4 we identify a suitable category of such regular regions in ${\mathbb{R}}^{n}$, the category ${\mathcal{C}\mathcal{S}}_{n}$ of fuzzy semilinear sets. We identify a Grothendieck topology on ${\mathcal{C}\mathcal{S}}_{n}$ and show that infinitesimal symmetries of any gapped state $\psi$ of a quantum lattice system on ${\mathbb{R}}^{n}$ can be described by a local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$ over ${\mathcal{C}\mathcal{S}}_{n}$. We also study gapped states invariant under an action of a compact Lie group $G$ and define a $G$-equivariant analog of ${\mathfrak{D}}^{\psi}_{al}$. In Section 6 we construct invariants of $G$-invariant gapped states. The construction is along the lines of [11, 8] and uses an inhomogeneous Maurer-Cartan equation. We show that these invariants are obstructions for promoting the symmetry $G$ to a local symmetry of the state $\psi$. We also explain how to generalize the construction of invariants to lattice systems defined on sufficiently nice (asymptotically conical) subsets of ${\mathbb{R}}^{n}$ and show that their invariants take values in a space which depends on the asymptotic geometry of the subset. This goes beyond what one can access using the TQFT heuristics. In Appendix A we isolate some proofs necessary for Section 3, and Appendix B develops some properties of the inhomogeneous Maurer-Cartan equation.
@@
2 Local Lie algebras
2.1 Locality and (pre-)cosheaves
As an example, consider the Lie algebra of gauge transformations, i.e. the Lie algebra ${\mathfrak{G}}(M):=C^{\infty}(M,{\mathfrak{g}})$ of smooth functions on $M$ with values in a finite-dimensional Lie algebra ${\mathfrak{g}}$. It is a global object attached to $M$. To “localize” it, for any open $U\subseteq M$ we define the Lie algebra ${\mathfrak{G}}(U)$ to be the space of smooth ${\mathfrak{g}}$-valued functions on $M$ whose closed support is contained in $U$. In particular, ${\mathfrak{G}}(\varnothing)=0$. For any inclusion of opens $U\subseteq V$ we have a homomorphism of Lie algebras $\iota_{VU}:{\mathfrak{G}}(U)\rightarrow{\mathfrak{G}}(V)$ such that the Lie algebras ${\mathfrak{G}}(U)$ assemble into a pre-cosheaf ${\mathfrak{G}}$ of Lie algebras on $M$. This is a coflasque pre-cosheaf, ie. all its structure maps $\iota_{VU}$ are injective. [^2] The terminology comes from sheaf theory, where a pre-sheaf ${\mathcal{F}}:Open(M)^{opp}\rightarrow{\mathcal{C}}$ is called flasque if for any $U\subseteq V$ the restriction morphism ${\mathcal{F}}(V)\rightarrow{\mathcal{F}}(U)$ is a surjection.
Here the first arrow is $\iota_{U,U\cap V}\oplus(-\iota_{V,U\cap V})$ and the second arrow is $\iota_{U\cup V,U}\oplus\iota_{U\cup V,V}$. Exactness follows from the existence of a partition of unity for the cover ${\mathfrak{U}}=\{U,V\}$ of $U\cup V$. In words, the exactness of the sequence (1) means that any element of ${\mathfrak{G}}(U\cup V)$ can be decomposed as a sum of elements of sub-algebras attached to $U$ and $V$ modulo ambiguities which take values in the sub-algebra attached to $U\cap V$.
More generally, the existence of a partition of unity implies that for any collection of opens $U_{i}$, $i\in I$, the following sequence is exact:
By definition, this means that the pre-cosheaf of vector spaces ${\mathfrak{G}}$ is a cosheaf of vector spaces. The cosheaf property is a compatibility of the pre-cosheaf with the notions of intersection and union of opens and expresses a stronger form of locality.
Note that the maps in the above exact sequences are not Lie algebra homomorphisms. Hence ${\mathfrak{G}}$ is not a cosheaf of Lie algebras. Nevertheless, the following additional property of ${\mathfrak{G}}$ can be regarded as a form of locality of the Lie bracket: for any $U,V,$ $[{\mathfrak{G}}(U),{\mathfrak{G}}(V)]\subseteq{\mathfrak{G}}(U\cap V)$. We will call this ”Property I”, where ”I” stands for ”intersection”. In particular, elements of ${\mathfrak{G}}(M)$ which are supported on non-intersecting opens $U$ and $V$ commute.
Symmetries of gapped states of quantum lattice spin systems are typically local only approximately. To phrase locality of symmetries in lattice systems in a similar language, one needs to replace the set $Open(M)$ of open subsets of $M$ with a more general structure which admits the notions of intersection, union, and cover.
For any pre-ordered set $(X,\leq)$ there is a natural notion of intersection and union. The intersection of $U,V\in X$ can be defined as the greatest lower bound (or meet) of both $U$ and $V$, i.e. a $W\in X$ such that $W\leq U$, $W\leq V$, and for any $W^{\prime}\leq U,V$ we have $W^{\prime}\leq W$. The meet of $U$ and $V$ is denoted $U\wedge V$. Similarly, the union of $U$ and $V$ can be defined as the smallest upper bound (or join) of both $U$ and $V$. It is denoted $U\vee V$. For a general pre-ordered set, the meet and join may not exist for all pairs of objects. If they exist, they are unique up to isomorphism. We will assume that $(X,\leq)$ is such that $U\wedge V$ and $U\vee V$ exist for all $U,V\in X$ [^3] If we turn $X$ into a poset by identifying isomorphic objects, then this means that the poset is a lattice in the sense of order theory. The existence of all pairwise meets and joins implies the existence of all finite meets and joins. In the case of the pre-ordered set $Open(M)$ arbitrary (i.e. not necessarily finite) joins make sense.
Finally, to define covers of elements of $X$, let us assume that $U\wedge(V\vee W)\leq(U\wedge V)\vee(U\wedge W)$ for all $U,V,W\in X$ [^4] The opposite relation is automatic, so this condition ensures that $U\wedge(V\vee W)\simeq(U\wedge V)\vee(U\wedge W)$ for all $U,V,W\in X$. This is equivalent to saying that the poset corresponding to $X$ is a distributive lattice. We will say that $X$ is a distributive pre-ordered set. This condition is certainly satisfied for $Open(M)$. We say that a collection ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ of elements of $X$ covers $A\in X$ iff $U_{i}\leq A$ for all $i\in I$ and $A\leq\bigvee_{i\in I}U_{i}$. This definition ensures that if ${\mathfrak{U}}$ covers $A$, then for any $B\leq A$ the collection ${\mathfrak{U}}\wedge B=\{U_{i}\wedge B\}_{i\in I}$ covers $B$.
Definition 2.1.
A local Lie algebra over $(X,\leq)$ is a coflasque pre-cosheaf of Lie algebras with Property I which is also a cosheaf of vector spaces over the corresponding site. A morphism of local Lie algebras is a morphism of the underlying pre-cosheaves of Lie algebras.
Remark 2.1.
Let ${\mathfrak{F}}$ be a local Lie algebra. Then for any $U\leq V$ the Lie algebra ${\mathfrak{F}}(U)$ is an ideal in ${\mathfrak{F}}(V)$. One can equivalently define a local Lie algebra as a pre-cosheaf of Lie algebras over $(X,\leq)$ which is a cosheaf of vector spaces and such that all co-restriction maps are inclusions of Lie ideals.
Remark 2.2.
In this paper all Lie algebras will be Fréchet-Lie algebras and all morphisms will be continuous. We define a local Fréchet-Lie algebra over $(X,\leq)$ to be a coflasque pre-cosheaf of Fréchet-Lie algebras with Property I which is also a cosheaf of vector spaces.
Lemma 2.1.
Let ${\mathfrak{F}}$ be a pre-cosheaf of vector spaces on a distributive lattice $X$, with co-restriction maps $\iota_{U,V}:U\to V$. Then ${\mathfrak{F}}$ is a cosheaf on $X$ (with the topology of finite covers) iff for any $U,V\in X$ the following sequence is exact

### v3

# A mathematical theory of topological invariants of quantum lattice systems

@@ Abstract
We show that Hall conductance and its non-abelian and higher-dimensional analogs are obstructions to promoting a symmetry of a state to a gauge symmetry. To do this, we define a local Lie system over a Grothendieck site as a pre-cosheaf of Lie algebras with additional properties and propose that a gauge symmetry should be described by such an object. We show that infinitesimal symmetries of a gapped state of a quantum lattice system form a local Lie system over a site of semilinear sets and use it to construct topological invariants of the state. Our construction applies to lattice systems on arbitrary asymptotically conical subsets of a Euclidean space including those which cannot be studied using field theory.
@@ 1 Introduction
1.1 Overview
A quantum lattice system is a mathematical model in quantum many-body physics in which degrees of freedom are arranged at points of a countable set $X\subset{\mathbb{R}}^{n}$ (“the lattice”) and interact locally. Such systems are naturally described in the language of operator algebras.
We begin by recalling three basic structures associated with a quantum lattice system [10, 11]. First of all, we are given a $C^{*}$-algebra $\mathscr{A}$ of quasi-local observables. This algebra admits local subalgebras $\mathscr{A}(Y)\subseteq\mathscr{A}$ for each subset $Y\subseteq X$. In the case of fermionic systems, $\mathscr{A}$ is ${\mathbb{Z}}_{2}$-graded. A state on $\mathscr{A}$ is a positive normalized linear functional encoding expectation values. A state is called gapped if it satisfies an appropriate energy stability condition. In this paper, we restrict attention to gapped states. Finally, one considers a Lie algebra of *-derivations of a dense *-subalgebra [^1] For infinite-dimensional $\mathscr{A}$, important derivations often do not extend to the entire algebra. of $\mathscr{A}$, whose elements are infinitesimal generators of dynamics.
For illustration, consider a bosonic system on a lattice consisting of a single point. Then $\mathscr{A}$ is a finite-dimensional simple $C^{*}$-algebra, and hence isomorphic to a matrix algebra $M_{n}(\mathbb{C})$. States correspond to density matrices, and gapped states are precisely vector states, which form the projective space $\mathbb{CP}^{n-1}$.
All derivations are inner and defined on the entire algebra, given by
$A\mapsto[F,A],$ (1)
for some $F\in M_{n}(\mathbb{C})$. Moreover, each derivation admits a unique traceless representative, so that the Lie algebra of derivations is isomorphic to $\mathfrak{sl}_{n}$.
Since the *-structure on $M_{n}(\mathbb{C})$ is given by the adjoint, a derivation is a *-derivation if and only if $F$ is skew-adjoint. Thus, the *-derivations correspond to the real Lie algebra $\mathfrak{su}_{n}$.
For a general lattice $\Lambda$, *-derivations need not be inner and may only be defined on a dense *-subalgebra of $\mathscr{A}$. Nevertheless, under suitable conditions, a *-derivation exponentiates to a one-parameter group of *-automorphisms of $\mathscr{A}$. More generally, a continuous path of *-derivations gives rise to a continuous path of *-automorphisms, which act on states by pullback and thus generate continuous families of states. This induces an equivalence relation: two states are considered equivalent if they are connected by such a path [26, 6]. A central problem is to construct invariants of gapped states under this relation. At present, there are satisfactory constructions only in one spatial dimension [17, 9]. Conjecturally, invariants of gapped states of lattice systems in higher dimensions are classified by Topological Quantum Field Theory (TQFT), but current tools seem inadequate to tackle this problem.
More can be said for the equivariant variant, which deals with gapped states invariant under a locality-preserving symmetry. A symmetry of a state by a compact connected Lie group $G$ can be described infinitesimally by a Lie algebra homomorphism into the *-derivations of $\mathscr{A}$ that preserve the state. Equivalently, the corresponding *-automorphisms fix the state. Such states are called $G$-invariant. This leads to a $G$-equivariant version of the problem: to construct invariants of $G$-invariant gapped states under continuous paths through $G$-invariant gapped states. When $\Lambda$ is a single point, such invariants correspond to one-dimensional representations of $G$. When $\Lambda$ is a countable set which uniformly fills ${\mathbb{R}}^{n}$ with $n>0$, invariants of $G$-invariant gapped states been recently constructed in [19] (see also [18, 3]). They take values in $G$-invariant polynomials on the Lie algebra of $G$. In the special case $G=U(1)$ and $n=2$ one obtains the famous Hall conductance [36].
1.2 Main Results and Outline
@@
In this paper, we develop a mathematical framework for defining invariants of gapped quantum lattice systems with Lie symmetries. In particular, we place the constructions of [18, 19, 3] on a more conceptual footing. A secondary goal is to generalize the construction in various directions. For example, we show how to define topological invariants of quantum lattice systems confined to well-behaved subsets of ${\mathbb{R}}^{n}$. This generalization makes explicit that the invariants take values in a vector space which is determined by the large-scale geometry of the subset.
The central idea of this paper is a new formulation of locality on a lattice, leading to the notion of a local Lie system. Building on the ideas introduced in [18, 19, 3], we associate to any (possibly unbounded) region of the lattice a space of derivations that are approximately localized on that region, and show that for sufficiently regular regions these spaces behave well under natural operations such as the commutator. These spaces assemble into a cosheaf on a suitable site, i.e. a category equipped with a Grothendieck topology, thereby producing a global geometric object encoding locality.
This structure provides a natural algebraic framework for symmetries. An on-site action of a Lie group gives rise to a representation by derivations, and we show that the invariants of [19] admit a purely algebraic description in this setting. Specifically, they arise as obstructions to promoting the symmetry to one acting by state-preserving derivations, with values in the homology of a differential graded Lie algebra (DGLA) naturally associated to the system. These obstruction classes pair with the Čech cohomology of the sphere at infinity to recover the invariants of [19]. This interpretation both clarifies their mathematical meaning and shows that they are independent of certain choices in their construction. It also highlights an analogy between topological invariants of gapped states and ’t Hooft anomalies in Quantum Field Theory which are obstructions for gauging a symmetry action, see e.g. [2] for a review.
@@
The content of the paper is as follows. In Section 2 we review quantum lattice systems and their symmetries. We introduce Lie algebras of derivations approximately localized on regions in ${\mathbb{R}}^{n}$ and prove some of their properties. In the next three sections we formalize these properties. In Section 3 we axiomatize the notion of an infinitesimal local symmetry by defining local Lie systems abstractly in terms of cosheaves on a site. We also introduce the Čech functor which assigns a DGLA to a local Lie system, and will serve as the main link between lattice geometry and algebraic topology. In Section 4 we identify a suitable category of regions in ${\mathbb{R}}^{n}$, the category ${\mathcal{C}\mathcal{S}}_{n}$ of fuzzy semilinear sets which comes with a natural Grothendieck topology. In Section 5 we show that infinitesimal symmetries of any gapped state $\psi$ of a quantum lattice system on ${\mathbb{R}}^{n}$ can be described by a local Lie system ${\mathfrak{D}}^{\psi}_{al}$ over ${\mathcal{C}\mathcal{S}}_{n}$. We also study gapped states invariant under an action of a compact Lie group $G$ and define a $G$-equivariant analog of ${\mathfrak{D}}^{\psi}_{al}$. In Section 6 we construct invariants of $G$-invariant gapped states. The construction is along the lines of [19, 3] and uses an inhomogeneous Maurer-Cartan equation. We show that these invariants are obstructions for promoting the symmetry $G$ to a local symmetry of the state $\psi$. We also explain how to generalize the construction of invariants to lattice systems defined on sufficiently nice (asymptotically conical) subsets of ${\mathbb{R}}^{n}$ and show that their invariants take values in a space which depends on the asymptotic geometry of the subset. This goes beyond what one can access using the TQFT heuristics. In Appendix A we isolate some proofs necessary for Section 2, and Appendix B develops some properties of the inhomogeneous Maurer-Cartan equation.
1.3 Related work
Recent progress has led to mathematically rigorous approaches to classifying gapped phases of quantum lattice systems, particularly in one and two dimensions [29, 28, 9, 17, 30, 35, 18, 3, 4, 5], and more generally in arbitrary dimension [19]. These works rigorously construct invariants of infinite-volume gapped states that are stable under suitable continuous deformations, and include well-known examples like the Hall conductance.
Acknowledgments
@@
2 Quantum lattice systems
This section is in part a review of quantum lattice systems on ${\mathbb{R}}^{n}$ following [10, 11] and [19]. We define the quasilocal algebra, its various subalgebras, and spaces of derivations. We then introduce spaces of derivations approximately localized on subsets of ${\mathbb{R}}^{n}$ and derive some of their properties. We also study spaces of derivations preserving a gapped state.
2.1 Observables and derivations
We use the $\ell^{\infty}$ metric on ${\mathbb{R}}^{n}$, i. e. $d(x,y):=\max_{i=1,\ldots,n}|x_{i}-y_{i}|$. For any $U,V\subseteq{\mathbb{R}}^{n}$ we write $\operatorname{diam}(U):=\sup_{x,y\in U}d(x,y)$ and $d(U,V):=\inf_{x\in U,y\in V}d(x,y)$. They take values in extended non-negative reals $0,\infty]$. Thus $\operatorname{diam}(\varnothing)=0$ and $d(U,\varnothing)=\infty$ for any $U\subseteq{\mathbb{R}}^{n}$. For a nonempty set $U$ and $r\geq 0$ we define $U^{r}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq r\}$ while we set $\varnothing^{r}=\varnothing$.
Definition 1.
A quantum lattice system consists of a countable subset $\Lambda\subseteq{\mathbb{R}}^{n}$ (“the lattice”) and a finite-dimensional ${\mathbb{Z}}_{2}$-graded complex Hilbert space $V_{x}$ for every $x\in\Lambda$ We assume that the number of points of $\Lambda$ in a hypercube of diameter $d$ is upper bounded by $C(1+d)^{n}$ for some $C>0$.  [^2] In [3] $\Lambda$ was taken to be ${\mathbb{Z}}^{n}$ for simplicity. In [19] $\Lambda$ was assumed to be a Delone set, i.e. it was required to be uniformly filling and uniformly discrete. These assumptions were imposed on physical grounds. All the results proved in [19] hold under weaker assumptions adopted in this paper.
Example 2.
Models describing fermionic particles have $V_{x}$ which is an exterior algebra of another vector space: $V_{x}=\wedge H_{x}$, where $H_{x}$ is a finite-dimensional Hilbert space (the space of one-particle states on site $x\in\Lambda$). The dimension of $H_{x}$ is referred to as the number of orbitals on site $x$. For translationally-invariant models, one typically takes all $H_{x}$ to be isomorphic to the same Hilbert space $H$, in which case $\dim H$ is called the number of bands.
For any $x\in\Lambda$ the algebra ${\mathscr{A}}_{x}=\mathrm{Hom}_{{\mathbb{C}}}(V_{x},V_{x})$ is a finite-dimensional ${\mathbb{Z}}_{2}$-graded $C^{*}$-algebra. For any nonempty bounded $X\subseteq{\mathbb{R}}^{n}$ let ${\mathscr{A}}(X):=\widehat{\bigotimes}_{x\in X\cap\Lambda}{\mathscr{A}}_{x}$, where $\widehat{\bigotimes}$ denotes super-tensor product of ${\mathbb{Z}}_{2}$-graded algebras. For any $X\subseteq Y$ there is an inclusion ${\mathscr{A}}(X)\hookrightarrow{\mathscr{A}}(Y)$ and the algebras ${\mathscr{A}}(X)$ form a direct system with respect to these inclusions. We extend this direct system to include the empty set by setting ${\mathscr{A}}(\varnothing)={\mathbb{C}}$ and letting the inclusion ${\mathscr{A}}(\varnothing)\hookrightarrow{\mathscr{A}}(X)$ take $\alpha\mapsto\alpha\boldsymbol{1}$. Each ${\mathscr{A}}(X)$ is a finite-dimensional ${\mathbb{Z}}_{2}$-graded $C^{*}$-algebra with the operator norm, and the inclusions ${\mathscr{A}}(X)\hookrightarrow{\mathscr{A}}(Y)$ preserve this norm. The normed ${\mathbb{Z}}_{2}$-graded *-algebra of local observables is